A sports car as rigid body needs to make a record jump from acertain ramp with some specific height covering certain horizontal distance.
1
Expert's answer
2020-12-23T06:55:18-0500


In order to make record, D should get max value. In order D to be max value, angle a should be "45^o" .

That is, H will be equal to L. And then, can say with formula:

"D=v_x \times t=v_0 \times cos(a) \times t;"

So then, if the car get "v_0" velocity at the end of the ramp and it will be t time in air and it jumps at a "45^o"angle, it will make record with D distance:

"D=v_0 \times cos(45^o) \times t=\frac{\sqrt2}{2} \times v_0 \times t;"